# Identifying Critical Dependencies in Large-Scale Continuous Software Engineering

### v1

@@ 2. Method
To identify the critical dependencies in large-scale CSE we conduct a qualitative survey relying on data from two large-scale Nordic fintech companies, SpareBank 1 Utvikling and NorEns, the pseudonyms are used to preserve the companies’ anonymity. Both companies have been focusing on the introduction of agile practices and methods for more than 10 years, promoting continuous improvement in software teams, and embraced hybrid software development (Conboy et al., 2023).
@@
Company Team ID Role (n. of interviews) Roles in the team Release frequency (approximately)
SB1U A1 TL (1), PO (1) Dev (4), TL (1) 4-8 times/day
A2 TL (2) Dev (3), TL (1), UX (1), PO (1), Team lead (1) 5-6 times/day
A3 TL (2) Dev (4), UX (1), PO (1) 4-5 times/day
A4 TL (1) Dev (2), TL (1) Several times/day
Gjensidige B1 TL (1), PO (1) BizDev (2), content designer (1), UX (1), Dev (5), TL (1), PO (1)
B2 TL (1), PO (1) Dev (6), UX (1), BizDev (1), business analyst (1), PO (1) 2-4 times/week or whenever ready
B3 TL (1), PO (1) Dev (7), TL (1), Content designer (1), UX (2), PO (1) 5-6 times/day
B5 TL (1), PO (1) PO (1), SM (1), UX (2), TL (1), Dev (5), DevOps (1) Bi-weekly or whenever ready
B6 TL (1), PO (1) Dev (6), UX (2), Content designer (1), PO (1) 5-6 times/day or whenever ready
@@
Note: TL = tech lead, PO = Product Owner, Dev = Developer, UX = UX-designer, BizDev = business developer, SM = Scrum Master.

### v3

@@ 2. Method
To identify the critical dependencies in large-scale CSE we conduct a qualitative survey relying on data from two large-scale Nordic fintech companies, SpareBank 1 Utvikling and Gjensidige. Both companies have been focusing on the introduction of agile practices and methods for more than 10 years, promoting continuous improvement in software teams, and embraced hybrid software development (Conboy et al., 2023).
@@
Company Team ID Role (n. of interviews) Release frequency in the team (approximately)
SB1U A1 TL (1), PO (1) 4-8 times/day
A2 TL (2) 5-6 times/day
A3 TL (2) 4-5 times/day
A4 TL (1) Several times/day
Gjensidige B1 TL (1), PO (1)
B2 TL (1), PO (1) 2-4 times/week or whenever ready
B3 TL (1), PO (1) 5-6 times/day
B5 TL (1), PO (1) Bi-weekly or whenever ready
B6 TL (1), PO (1) 5-6 times/day or whenever ready